OpenStudy (alones.):

Hrmm I can't figure it out :3

OpenStudy (alones.):

12 a^3p^4 ÷ -2a^2p i got −6a^5p^5 and -6ap^3

OpenStudy (faiqraees):

-6ap^3 is correct

OpenStudy (faiqraees):

Remember when there is a division sign, the powers have to be subtracted 

OpenStudy (faiqraees):

oh sorry it should be -2 instead of -6 in the diagram

OpenStudy (alones.):

Wait so I was right last time :3

OpenStudy (alones.):

Wait so what is it??

OpenStudy (faiqraees):

-6ap^3, this is correct the other one is wrong

OpenStudy (alones.):

@CandyCove Is he right.. cuz i keep on gettin the other one -.-

OpenStudy (faiqraees):

I am correct :(

candycove (candycove):

I don't know any of this. :))

OpenStudy (alones.):

Cove -.-

candycove (candycove):

Erica -.-

OpenStudy (alones.):

@imqwerty -.-

imqwerty (imqwerty):

\(\Large \frac{12 a^3p^4}{ -2a^2p}\) remember this rule and try applying it-> \(\Large \color{red}{\frac{x^p}{x^q}=x^{p-q}}\) now tell your final answer

candycove (candycove):

Smarty pants. :))

OpenStudy (alones.):

K

OpenStudy (alones.):

LOL COVE why pants not head??

candycove (candycove):

It's a term, Erica. -.-

OpenStudy (alones.):

same :3

imqwerty (imqwerty):

:) (:

imqwerty (imqwerty):

just seeing what it wuld luk like \(\Huge \color{brown}{ 💩}\)

OpenStudy (alones.):

Qwerty really -.-

imqwerty (imqwerty):

LOL!!!

OpenStudy (alones.):

you make it worse -.-cove qwerty so there right :3

candycove (candycove):

\(\Huge\color{black}{👊}\)

imqwerty (imqwerty):

what is that @candycove :o ?

OpenStudy (alones.):

it's a fist in the face..i wonder how did she make it..her face is turned away but her fist is turning my way -.-

candycove (candycove):

HUEHUEHEUHEUHEEUHEUEH

imqwerty (imqwerty):

lol Your way!!! xD

OpenStudy (alones.):

I mean in my face

OpenStudy (alones.):

qwerty so it's right ?

imqwerty (imqwerty):

candy wants to beat you? :o

candycove (candycove):

\(\Huge\color{teal}{💙}\)

OpenStudy (alones.):

No she gave me a fist five

OpenStudy (alones.):

**four

candycove (candycove):

A <3 for Erica.

OpenStudy (*poker-face*):

see \(\Huge \color{red}{❤}\) like a rose petal :*

OpenStudy (alones.):

GEEZ IT'S NOT LaTex -.- i need to know if i'm right GAWD -.-

OpenStudy (alones.):

GAWD DAANG!=.-

candycove (candycove):

ERICAAAA \(\Huge\color{black}{👻}\)

OpenStudy (alones.):

@pooja195 Do i look like a ghost?!? FEEZ I'M LOOSING MY PATIANCE

OpenStudy (*poker-face*):

wow qwerty thats so good!! (':

OpenStudy (alones.):

Wow qwerty o lala this took a while for you to draw

pooja195 (pooja195):

Stop spamming

imqwerty (imqwerty):

HA!!!! candycove got an eye shut

candycove (candycove):

\(\Huge\color{purple}{Pooja!}\)

OpenStudy (alones.):

-.-

OpenStudy (alones.):

Pooja did i get it right or that dude was correct

OpenStudy (*poker-face*):

pooja can you help me?

candycove (candycove):

\(\Huge\color{black}{-_--}\)

imqwerty (imqwerty):

LOL that LOOKS lke a fuglleyy frogggg

imqwerty (imqwerty):

alone did you get the answer tho?

candycove (candycove):

\(\Huge\color{black}{:(}\)

OpenStudy (alones.):

Thanks qwerty for remebering that I'm still here looking at your lazy drawings...i did I'm asking if I got it right -.-

imqwerty (imqwerty):

\(\Huge \ddot \frown\)

imqwerty (imqwerty):

yeah so whats the answer? :0

imqwerty (imqwerty):

* :)

OpenStudy (alones.):

YOU STILL Don't know the answer -.- i got 2 of them look in the top -.-

OpenStudy (alones.):

Geez my brain is hot i have to put it in the refridge. DID I GET IT RIGHT!!!?

OpenStudy (*poker-face*):

use the identity that qwert gave

OpenStudy (alones.):

I did and i got the same answer -.-

imqwerty (imqwerty):

what did you do?

OpenStudy (alones.):

i didn't do anything -.-

imqwerty (imqwerty):

alright i'll go step wise and you tell if you get stuck anywhere :)

OpenStudy (alones.):

i keep on getting this one −6a^5p^5

imqwerty (imqwerty):

we got this- \(\huge\frac{12 a^{\color{violet}3}p^{\color{violet}4}}{ -2a^{\color{violet}2}p^{\color{violet}1}}\) now i use this identity-\(\Large \color{red}{\frac{x^p}{x^q}=x^{p-q}}\) and we get this- \(\huge\frac{12 a^{\color{violet}{3-2}} p^{\color{violet}{4-1}}}{ -2}\)

imqwerty (imqwerty):

alright till here?

OpenStudy (alones.):

yes

imqwerty (imqwerty):

okay try to simplify it from here :))

OpenStudy (alones.):

K i got 12a^1^3

imqwerty (imqwerty):

what about that -2 down there in the denominator? :)

OpenStudy (alones.):

12

imqwerty (imqwerty):

noo well check if that \(-2\) divides the 12 and if something cancels out \(\huge \frac{\color{red}{12} a^1 p^3}{\color{red}{-2}}\)

OpenStudy (alones.):

Oh xP my bad 5

OpenStudy (alones.):

**6

imqwerty (imqwerty):

yes but don't you that that will be negative 6 :) ?

OpenStudy (alones.):

oh yea so it looks like -6a1p3

imqwerty (imqwerty):

yes correct :)

OpenStudy (alones.):

wait so it's that O.O

imqwerty (imqwerty):

yess :) we only had to simplify

OpenStudy (alones.):

Oh dumb dumb.. then mine was qrong..k thanks<3

imqwerty (imqwerty):

np :)